OpenStudy (anonymous):

i have no clue how to set up a formula for this one: Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and Train B is traveling at 100 miles per hour. Train A passes a station at 5:15 pm. If Train B passes the same station at 5:30 pm, at what time will Train B catch up to Train A?

OpenStudy (anonymous):

d = r * t Train A d = 80(t + .25) Train A basically gets a 15 minute head start.. but 15 minutes in hours is 1/4 of an hour which is .25 because time must be in hours. Train B = d = 100(t) Because they are traveling on the same track their distances will be equal to each other. They both actually start at the station and when they catch each other they have travelled the same distance

OpenStudy (anonymous):

80(t + .25) = 100t 80t + 20 = 100t 20 = 20t t = 1 It will take Train B 1 hour to catch Train A

OpenStudy (anonymous):

so Train B will catch up with Train A at 6:30?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok great, thx!